I would appreciate any help you can give me.
A spherical water tank is 12.6 ft. in diameter. What is the weight of the water in the tank if water weighs 62.4 lb/ft cubed?
Thank You,
makennedy01

Answer by stanbon(75887)

A spherical water tank is 12.6 ft. in diameter. What is the weight of the water in the tank if water weighs 62.4 lb/ft cubed?
---------------
Volume = (4/3)(pi)(6.3)^3
Volume = 1047.39 cu ft
Weight = (1047.39)*62.4 = 65357.41 lbs.
------------
Cheers,
Stan H.